"E4-22A Using ABC to bill clients at a service firm (Learning Objective 2) Kushner & Company is an architectural firm specializing in home remodeling for private clients and new office buildings for corporate clients. Kushner charges customers at a billing rate equal to 130% of the client’s total job cost. A client’s total job cost is a combination of (1) professional time spent on the client ($63 per hour cost of employing each professional) and (2) operating overhead allocated to the client’s job. Kushner allocates operating overhead to jobs based on professional hours spent on the job. Kushner estimates its five professionals will incur a total of 10,000 professional hours working on client jobs during the year. All operating costs other than professional salaries (travel reimbursements, copy costs, secretarial salaries, office lease, and so forth) can be assigned to the three activities. Total activity costs, cost drivers, and total usage of those cost drivers are estimated as follows: Activity Total Activity Cost Cost Driver Total Usage by Corporate Clients Total Usage by Private Clients Transportation to clients $ 6,000 Round-trip mileage to clients 4,500 miles 10,500 miles Blueprint copying 32,000 Number of copies 250 copies 750 copies Office support 195,000 Secretarial time 2,500 secretarial hours 2,500 secretarial hours Total operating overhead $ 233,000 Yimeng Li hired Kushner to design her kitchen remodeling. A total of 25 professional hours were incurred on this job. In addition, Li’s remodeling job required one of the professionals to travel back and forth to her house for a total of 125 miles. The blueprints had to be copied four times because Li changed the plans several times. In addition, 15 hours of secretarial time were used lining up the subcontractors for the job.

Requirements Calculate the current indirect cost allocation rate per professional hour.

Calculate the total amount that would be billed to Li given the current costing structure.

Calculate the activity cost allocation rates that could be used to allocate operating overhead costs to client jobs.

Calculate the amount that would be billed to Li using ABC costing.

Which type of billing system is more fair to clients? Explain."
Answer #1

Part 1

Operating overhead

Total professional hours

Current operating overhead allocation rate

$233,000

÷

10000

=

$23.30 per professional hour

Part 2

Total cost of job = ($63*25)+($23.30*25) = $2157.50

Amount billed = $2157.50 * 130% = $2804.75

Part 3

Activity

Cost

Total activity allocation base

Activity allocation rate

Transportation to clients

6000

÷

15,000 miles driven

=

$0.40 per mile

Blueprint copying

32000

÷

1,000 copies

=

$32 per copy

Office support

195000

÷

5000 secretarial hours

=

$39 per secretarial hour

4500+10500 = 15000

250+750 = 1000

2500+2500 = 5000

Part 4

Total cost job = (63*25)+(125*0.4)+(4*32)+(15*39) = 2338

Amount billed = Total cost job*markup percentage = 2338*130% = $3039.40

Part 5

The ABC billing system is considered to be more fair to clients as in this system the costs are charged according to the usage of resources. Taking example copying blueprints, this cost seems to be very expensive in traditional billing system. However, according to ABC system, this cost will be charged to clients as per the number of blueprint copies that their job required. In ABC system, the use of cost driver is done to recognize the operating costs that are incurred by each unique client job.